OpenStudy (anonymous):

Help PLEASE I really don't understand this, √2(√8+4√3) thank you(:

OpenStudy (anonymous):

is the answer simplified ?

OpenStudy (anonymous):

I don't have an answer for that

OpenStudy (anonymous):

give me ur anwser choices

OpenStudy (anonymous):

*answer

OpenStudy (anonymous):

okay A.√10+4√6 B.4+4√6 C.2√2+4√6

OpenStudy (anonymous):

http://www.freemathhelp.com/equation-solver.php This is a great calculator for ANY equation. Just thought I'd share what has been very useful to me. =)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

you got it ??

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

its b!

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

your welcome (:

OpenStudy (anonymous):

(: ive been really sick so im behind and i dont understand all of this

jimthompson5910 (jim_thompson5910):

Distribute, then combine the roots and simplify if possible \[\large \sqrt{2}(\sqrt{8}+4\sqrt{3})\] \[\large \sqrt{2}*\sqrt{8}+\sqrt{2}*4\sqrt{3}\] \[\large \sqrt{2}*\sqrt{8}+4*\sqrt{2}\sqrt{3}\] \[\large \sqrt{2*8}+4*\sqrt{2*3}\] \[\large \sqrt{16}+4*\sqrt{6}\] \[\large 4+4\sqrt{6}\]

jimthompson5910 (jim_thompson5910):

so this means \[\large \sqrt{2}(\sqrt{8}+4\sqrt{3}) = 4+4\sqrt{6}\]

OpenStudy (anonymous):

lol its ok im in like a similar situation

OpenStudy (anonymous):

Thanks for all the help yah its hard